roll a pair of dice and let I=1 if the sum is 7 and I=0 if sum is others. find the variance of I.
Answer #1

Total outcomes when rolling a pair of dice = 36

Total ways to get sum 7= 6

P(7) = 6/36 = 1/6

So, P(I = 1) = 1/6

P(I = 0) = 1 - 1/6 = 5/6

I 0 1 Total
P 5/6 1/6 1
I*P 0 1/6 1/6
I^2 *P 0 1/6 1/6

Hence 5/36 is the Variance of I.

Please comment if any doubt. Thank you.